Given the following information, what is the Total Annual Cost when ordering and holding costs only are taken into consideration?

Annual Demand 15,000
Order Quantity 500
Order Cost $30 per order
Holding Cost 20% per unit per year
Unit Cost                                 25
Answer #1

DEMAND = 15000
ORDERING COST = 30
HOLDING COST = 5
COST PER UNIT = 25

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST = (500 / 2) * 5 = 1250
ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST = (15000 / 500) * 30 = 900
TOTAL COST OF MANAGING = AHC + AOC = 1250 + 900 = 2150